the susceptibility of ferromagnetic materials is K at 27o C.at what temperature will its susceptibility be 0.5K___ 

Grade:12

1 Answers

Sambhav Jain
6 years ago
Susceptibility of ferromagnetic material
Xm=C/T
Given Susceptibility of ferromagnetic material at 27 degree Celcius is
Xm=K
Xm1=0.5k
T=27+273=300k
Xm1/Xm=T/T1
T1=Xm* T/Xm1=300 * K/0.5K=600K